Assume that the volume of a sphere is numerically equal to twice its surface area Find the radius of the sphere
Answer by srinivas.g(540)

volume of sphere = (4/3 )Pi r^ 3
surface area = 4 pi r^2
volume = 2 *surface area
(4/3 ) pi r^3 = 2 (4 pi r^2 )
r=6 m